OpenStudy (bobobox):

The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger? (2 points) 42 56 182 194

OpenStudy (bobobox):

@AlexandervonHumboldt2

OpenStudy (anonymous):

Let's say \(x\) is number of adult visitors (18 and older) and \(y\) of young visitors (18 or younger). We know that in total they make 224. so we say: $$ x + y = 224 \implies x = 224 - y $$We also know that every adult visitor paid $12 and every young visitor paid $8. In total they all paid $2,520. So we say: $$ 12x + 8y = 2520 $$But as shown above, we can describe 'x' as (total_visitors - young_visitors). So: $$ 12(224-y) + 8y = 2520 \\ 2688 - 12y + 8y = 2520 \\ -4y = 2520 - 2688 \\ -4y = -168 \\ y = \frac{168}{4} = 42 $$And we found \(y\), which is the number of the young visitors. Hope it's clear.